OpenStudy (anonymous):

sin165 (find exact value by using sum or difference identity)

OpenStudy (raden):

hint : 165 = 210-45

OpenStudy (anonymous):

haha good point

OpenStudy (anonymous):

once you get that what do you do after that?

OpenStudy (raden):

use the formula : sin(a-b) = sincosb - cosasinb

OpenStudy (anonymous):

right! :)

OpenStudy (anonymous):

I have another one..

OpenStudy (anonymous):

tan285

OpenStudy (raden):

285=240+45 use the formula: tan(a+b) = (tana + tanb)/(1-tanatanb)

OpenStudy (anonymous):

can you get then answer while i get it so I can make sure it is right

OpenStudy (raden):

hmmmm... u first use, tan240=sqrt(3), tan45=1 tell m what u get after that

OpenStudy (anonymous):

I got sqrt(3)+1 over 1-sqrt(3)

OpenStudy (raden):

yes, correct actually, u need rationalize of it right ? simplify it

OpenStudy (anonymous):

it would equal to one right

OpenStudy (raden):

a fraction would be 1 if numerator and denominator was same

OpenStudy (anonymous):

oh when rationalized i got -2

OpenStudy (raden):

multiply by conyugate of denominator that fraction

OpenStudy (anonymous):

NM

OpenStudy (raden):

it means